# with Hierarchical Adaptive Screening

### v1

@@ 1 Introduction
  • Upon the triggering of IRS indicator, we develop a Hierarchical Adaptive Screening (HAC) that adaptively balances efficiency and accuracy. Supported by four stage-specific attack vector databases across the agent lifecycle (Query, Planning, Action, Observation), HAC resolves known threats via fast vector matching while routing unfamiliar cases to a deep-reasoning path for autonomous adjudication.
@@ 3.2 Overview
The Spider-Sense framework enhances the security of autonomous agents to address inherent vulnerabilities. Our approach enables the agent to dynamically and autonomously sense potential risks throughout its execution; for any identified risk, the system invokes the Hierarchical Adaptive Screening (HAC) (Section 3.4) to perform a security inspection. We term this endogenous capability, which endows the model with self-driven defense, as Intrinsic Risk Sensing (IRS) (Section 3.3).
@@
Upon sensing a potential threat, the agent deterministically wraps the suspicious artifact in a specialized template, which is then routed to the Hierarchical Adaptive Screening (HAC) mechanism for hierarchical verification of the detected risk. By balancing efficient pattern matching (coarse-grained detection) with deep, deliberative reasoning (fine-grained analysis), HAC enables adaptive threat validation, after which the main agent autonomously decides whether to continue or terminate execution, without compromising operational efficiency.
@@ 3.4 Hierarchical Adaptive Screening
IRS decides when to perform inspections via $\phi_{t}^{(k)}$ and extracts a stage artifact $p_{t}^{(k)}$ using the corresponding template. Once triggered, $p_{t}^{(k)}$ is routed to a stage-specific inspector, implemented as a Hierarchical Adaptive Screening (HAC). HAC combines fast Coarse-grained Detection with slower, more in-depth Fine-grained Analysis in a hierarchical manner to enable adaptive inspection scheduling. Specifically, lightweight screening is applied when the fast detection has high confidence, while stronger and more time-consuming inspection stages are triggered as confidence decreases. This design dynamically adjusts the timing and intensity of inspection and ultimately returns a concise checking result to the agent.
To enable such fast detection, HAC maintains stage-wise attack vector databases to support retrieval-based defense, with construction details provided in Appendix D. For each stage $k\in\mathcal{K}$, we construct a case bank $\mathcal{D}^{(k)}$ that stores vectorized representations of commonly observed attack patterns from existing datasets.
@@
Based on the stage-wise vector database, HAC performs Coarse-grained Detection by measuring the similarity between the current artifact and stored attack patterns using cosine similarity. Given a stage artifact $p_{t}^{(k)}$, the corresponding inspector embeds it as a vector representation $\mathbf{v}_{t}^{(k)}$ and computes its similarity to each case $\mathbf{v}_{i}^{(k)}\in\mathcal{D}^{(k)}$ as
@@
Otherwise, when the similarity score falls below the threshold, i.e., $s_{t}^{(k)}<\tau^{(k)}$, HAC escalates to a Fine-grained Analysis stage that invokes more deliberative reasoning. In this case, the inspector retrieves the top-$K$ most similar cases from the stage-wise case bank $\mathcal{D}^{(k)}$ based on cosine similarity,
@@ 5 Experiments
This section presents a comprehensive experimental evaluation of Spider-Sense. Section 5.1 describes the experimental setup, including the benchmarks, baseline methods, and evaluation metrics. Section 5.2 reports the main experimental results. Section 5.3 provides ablation studies on the proposed IRS and HAC components. Finally, Section 5.4 presents a representative case study demonstrating agent execution under Spider-Sense. Additional experimental details are provided in Appendix A.

### v2

@@ 1 Introduction
  • Upon the triggering of IRS indicator, we develop a Hierarchical Adaptive Screening (HAS) that adaptively balances efficiency and accuracy. Supported by four stage-specific attack vector databases across the agent lifecycle (Query, Planning, Action, Observation), HAS resolves known threats via fast vector matching while routing unfamiliar cases to a deep-reasoning path for autonomous adjudication.
@@ 3.2 Overview
The Spider-Sense framework enhances the security of autonomous agents to address inherent vulnerabilities. Our approach enables the agent to dynamically and autonomously sense potential risks throughout its execution; for any identified risk, the system invokes the Hierarchical Adaptive Screening (HAS) (Section 3.4) to perform a security inspection. We term this endogenous capability, which endows the model with self-driven defense, as Intrinsic Risk Sensing (IRS) (Section 3.3).
@@
Upon sensing a potential threat, the agent deterministically wraps the suspicious artifact in a specialized template, which is then routed to the Hierarchical Adaptive Screening (HAS) mechanism for hierarchical verification of the detected risk. By balancing efficient pattern matching (coarse-grained detection) with deep, deliberative reasoning (fine-grained analysis), HAS enables adaptive threat validation, after which the main agent autonomously decides whether to continue or terminate execution, without compromising operational efficiency.
@@ 3.4 Hierarchical Adaptive Screening
IRS decides when to perform inspections via $\phi_{t}^{(k)}$ and extracts a stage artifact $p_{t}^{(k)}$ using the corresponding template. Once triggered, $p_{t}^{(k)}$ is routed to a stage-specific inspector, implemented as a Hierarchical Adaptive Screening (HAS). HAS combines fast Coarse-grained Detection with slower, more in-depth Fine-grained Analysis in a hierarchical manner to enable adaptive inspection scheduling. Specifically, lightweight screening is applied when the fast detection has high confidence, while stronger and more time-consuming inspection stages are triggered as confidence decreases. This design dynamically adjusts the timing and intensity of inspection and ultimately returns a concise checking result to the agent.
To enable such fast detection, HAS maintains stage-wise attack vector databases to support retrieval-based defense, with construction details provided in Appendix D. For each stage $k\in\mathcal{K}$, we construct a case bank $\mathcal{D}^{(k)}$ that stores vectorized representations of commonly observed attack patterns from existing datasets.
@@
Based on the stage-wise vector database, HAS performs Coarse-grained Detection by measuring the similarity between the current artifact and stored attack patterns using cosine similarity. Given a stage artifact $p_{t}^{(k)}$, the corresponding inspector embeds it as a vector representation $\mathbf{v}_{t}^{(k)}$ and computes its similarity to each case $\mathbf{v}_{i}^{(k)}\in\mathcal{D}^{(k)}$ as
@@
Otherwise, when the similarity score falls below the threshold, i.e., $s_{t}^{(k)}<\tau^{(k)}$, HAS escalates to a Fine-grained Analysis stage that invokes more deliberative reasoning. In this case, the inspector retrieves the top-$K$ most similar cases from the stage-wise case bank $\mathcal{D}^{(k)}$ based on cosine similarity,
@@ 5 Experiments
This section presents a comprehensive experimental evaluation of Spider-Sense. Section 5.1 describes the experimental setup, including the benchmarks, baseline methods, and evaluation metrics. Section 5.2 reports the main experimental results. Section 5.3 provides ablation studies on the proposed IRS and HAS components. Finally, Section 5.4 presents a representative case study demonstrating agent execution under Spider-Sense. Additional experimental details are provided in Appendix A.